Hi - my tutor is out of town and am stuck on this problem. Solve this sytem of equations for real number solutions: y=x^2-4 and x^2 = y^2 = 10 So far I have done this:
x^2 + y^2 = 10
-X^2 = y = -4
y^2 + y = 6
y^2 + y - 6 = 0
(y-2) (y + 3) = 0
y-2 = 0 y + 3 = 0
Y = 2 and y = -3
am now stuck
Thanks for your assistance
Carol

Answer by Alan3354(69443)

y=x^2-4 and
x^2 = y^2 = 10
--------------
So far I have done this:
x^2 + y^2 = 10
-X^2 = y = -4
y^2 + y = 6
y^2 + y - 6 = 0
(y-2) (y + 3) = 0
y-2 = 0 y + 3 = 0
Y = 2 and y = -3
am now stuck
------------
y=x^2-4 and
x^2 + y^2 = 10
Sub for y
x^2 + (x^2-4)^2 = 10
x^4 -7x^2 + 6 = 0
(x^2 - 6)*(x^2 - 1) = 0
x = ± sqrt(6) ; y = 2
x = ± 1 ; y = -3